Train A leaves a station traveling at 60 miles per hour. Six hours later train B leaves the same station traveling in the same direction at 80 miles per hour. How long does it take for train B to catch up to train A. Please Help Me!!
Answer by edjones(7311)

d=st d=dist s=speed t=time
train A's d: 60(t+6)
train B's d: 80t
60(t+6)=80t
60t+360=80t
60t-60t+360=80t-60t
20t=360
20t/20=360/20
t=18 hrs
.
Check:
60*24=1440
80*18=1440
.
Ed